The pH of pure water at 60°C is 8.517. Which one of the following statements is false ? The pOH of pure water is 8.517 at 60°C pKw 17.03 at 60°C Pure water is basic at 60°C Ku = 9.25 10-18 at 60°C. Pure water has a greater percent ionization at 60°C than at 25°C.
Answer #1

Solution:- If we consider the ionization of water then it gives H+ and OH- and the concentrations are equal.

Given pH is 8.571

So, [H+] = 10-8.571 = 2.685 x 10-9 M

Since, [H+] = [OH-]

So, [OH-] = 2.685 x 10-9 M

pOH = - log 2.685 x 10-9

pOH = 8.571

For water,

pKw = pH + pOH

pKw = 8.517 + 8.517

pKw = 17.03

We know that, Kw = 10-pKw

Kw = 10-17.03

Kw = 9.33 x 10-18

This calculated value is almost equal to the given value of Kw so This one is True.

Kw value increases as the temperature is increased and so the percent ionization also increases.

Pure water has equal concentrations of H+ and OH- that makes it neutral. So, the false statement is, " pure water is basic at 60 degree C."